At one point in the Illinois state lottery Lotto game, a person was required to choose six numbers (in any order) among 44 numbers. In how many ways can this be done?


1
Expert's answer
2021-11-05T13:48:42-0400

"n=C^6_{44}=\frac{44!}{38!6!}=7059052"